### v1

# Real-Time and Accurate: Zero-shot High-Fidelity Singing Voice Conversion with Multi-Condition Flow Synthesis

Abstract
Singing voice conversion is to convert the source sing voice into the target sing voice except for the content. Currently, flow-based models can complete the task of voice conversion, but they struggle to effectively extract latent variables in the more rhythmically rich and emotionally expressive task of singing voice conversion, while also facing issues with low efficiency in speech processing. In this paper, we propose a high-fidelity flow-based model based on multi-decoupling feature constraints, which enhances the capture of vocal details by integrating multiple encoders. We also use iSTFT to enhance the speed of speech processing by replacing some layers of the Vocoder. We compare the synthesized singing voice with other models from multiple dimensions, and our proposed model is highly consistent with the current state-of-the-art, with the demo which is available at https://lazycat1119.github.io/RASVC-demo/
Index Terms:
Deep learning, Singing voice conversion, iSTFT, Flow model
I Introduction
The main task of voice conversion (SVC) is to convert the voice of the source singer into the voice of the target speaker, while retaining the original lyrics and melody. High-quality singing conversion not only needs to accurately retain the content information but also needs to capture and reconstruct the information of the singer’s musical elements (including timbre, pitch, rhythm, emotion, etc.)[1][2]
Speech content should fully contain all the content information. The quality of PPGS-based representation [3][4][5][6]proposed by early researchers depends heavily on the performance of the ASR system, which requires a lot of labeled data for training. Later, SSL-based representation[7]can decouple the speaker’s information, but its generalization ability and stability are poor. VAE-based content embedding is stable and can capture deep information at the same time[8][9][10], but it still entangles speaker information, such as WavLM[11] and ContentVec[12][13]. These models all take continuous variables as the existing form of content information, and now the proposed speech discretization accelerates the speed of speech processing, such as VQ-VAE[14][15] and RVQ[16][17], but they are at the expense of losing some content information.
On the generation framework, one kind of model adopts an end-to-end system, such as FastSpeech[18], AutoVC[19], or Transformer TTS[20][21], and the other kind of model generates acoustic features (such as Mel spectrogram). It uses a vocoder to generate speech waveforms. At present, the vocoder based on GANs is hot, and high-fidelity waveforms, such as HiFi-GAN[22][23] and MelGAN[24], are generated through confrontation training, but the problem is that the decoding speed is too slow. In recent years, the vocoder model based on Flow has also been widely concerned[25][26]. It is a generation model that makes the optimization goal solvable by constructing reversible transformations, such as Nice[27], RealNVP[28][29], Glow[30], and WaveGlow[31]. They accelerate the process of speech generation through parallelization. However, in these speech models, the flow input conditions are few, and only the speaker’s identity embedding and content embedding are considered, which leads to the lack of naturalness and expressiveness of the generated speech and makes it difficult to perform the task of song conversion.
To solve the above problems, we use the HuBERT-Soft model to complete content extraction, which not only reduces the dependence on a large number of labeled data but also directly finds an accurate balance between discrete features and continuous features and separates content information[32]. To increase the naturalness of singing, we use the flow model to generate the distribution of potential variables reversibly, and at the same time, we introduce timbre encoder, pitch encoder, and emotion encoder as conditions, which improves the information integrity of the generated speech. Finally, we choose a simple inverse short-time Fourier transform (iSTFT) to skip some complicated vocoder processing steps, speed up the processing of the decoder module, and improve the reasoning speed of the model.
Figure 1: Overview of model structure
In this study, we have realized a fast and high-quality singing conversion model. The main contributions of this paper can be summarized as follows:
1. We propose to use the HuBERT-Soft model and combine the VITS model with HuBERT-Soft, a feature extraction model of voice content, to effectively extract the content information in singing.
2. Introduce and improve the flow model, not only to extract the speaker’s timbre, but also to introduce pitch and emotion as the condition of the flow by using the f0 extractor and emotion extractor, and to generate new variables by using the reversibility of the flow, which greatly improves the naturalness and expressiveness of singing conversion.
3. To speed up the VITS decoder module, we choose a simple inverse short-time Fourier transform (iSTFT) to skip some complicated vocoder processing steps and directly convert from frequency domain features to time domain waveforms.
II Method
II-A Overall pipeline
As shown in Figure 1, the entire model includes an F0 encoder, an emotion encoder, a speaker encoder, a content encoder, a Multi-condition flow, an MS-iSTFT-Decoder, and a discriminator, among which only the discriminator follows the VITS model[33]. In the following text, we focus on describing the aforementioned multiple encoders, the Multi-condition flow, and the MS-iSTFT-Decoder.
II-B Multi-condition encoder
In this section, we will introduce four encoders used to assist speech generation before entering the flow-based model, namely content Encoder, speaker encoder, f0 encoder, and emotion encoder.
II-B1 Content Encoder
Content encoding includes the Hubert-soft model and a content encoder. A major challenge in self-supervised speech representation learning is that speech contains multiple units and there are typically no discrete words or characters as input. By inputting the source speech into the pre-trained Hubert-soft model, we learn the latent distribution of the discrete variables of speech, which results in an aligned sequence $h(x)=Z=[z_{1},\ldots,z_{T}]$, where each discrete unit has 1024-dimensional features. These features are then fed into the content encoder and transformed into a lower-dimensional content embedding $Z_{c}$.
II-B2 Speaker Encoder
The Mel feature of the speaker’s voice can best express the different identities of the speaker[34], so we input the Mel feature of the target singer into a simple LSTM framework, and we believe that the speaker encoder will output the identity information about the speaker.
II-B3 F0 Encoder
We employ a monophonic pitch tracker based on a deep convolutional neural network to obtain the continuous fundamental frequency (F0). Considering the differences in fundamental frequencies between male and female voices can lead to unnatural-sounding singing, during the inference phase, we add the average difference in fundamental frequencies between the target and source voices to the source singing’s F0. This results in a final F0 feature that not only preserves the dynamic characteristics of the original song’s vocal track but also retains the vocal traits of the target speaker.
II-B4 Emotion Encoder
We use a self-supervised and pre-trained universal emotional expression model, emotion2vec[35], to extract the emotional features of each song, serving as a constraint for the flow model, which is beneficial for a more natural transition in song conversion and the embodiment of the song’s artistic conception.
II-C Flow with multiple decoupling attribute constraints
Similar to VITS, we introduce a standardized stream composed of multiple coupling affine layers. The coupling layers are reversible throughout the entire network and conserve volume as 1 in Jacobian form. Within the coupling layers, we use the decoupled features output by multiple encoders as constraints, obtaining a more refined latent distribution $Z$ by inputting the content embedding.
TABLE I: Model comparison
Method MOS/Naturalness MOS/Similarity PESQ Voice similarity RTF(GPU)
FastSVC 3.52 ± 0.10 3.27 ± 0.22 - 0.433 0.031
SoVITS-Flow 3.10 ± 0.22 2.90 ± 0.23 2.486 0.585 0.008
CoMoSVC $\mathbf{4.27\pm 0.16}$ 4.00 ± 0.19 $\mathbf{2.948}$ 0.585 $\mathbf{0.006}$
DiffSVC 4.23 ± 0.19 3.95 ± 0.21 2.917 0.598 0.278
Multi-F0 Model 3.85 ± 0.02 3.47 ± 0.04 - - -
UCD-SVC 3.06 ± 0.08 2.67 ± 0.18 - 0.588 0.103
$\mathbf{Ours}$ 4.14 ± 0.17 $\downarrow$ $\mathbf{4.02\pm 0.19}$$\uparrow$ 2.834 $\downarrow$ $\mathbf{0.603}$uparrow 0.048 $\downarrow$
Compared with the indexes of the restored songs at 16kHz, many experiments have been done
to get the average value, and bold indicates the best result.
II-D MS-iSTFT-Decoder
We utilize iSTFT to replace certain repetitive network layers in the previous HiFi-GAN vocoder by introducing the computation of phase and amplitude, converting discrete frequency domain representations (such as Mel spectrograms) into continuous time-domain waveforms. This approach effectively reduces computational load and accelerates the audio synthesis process.
II-E Multi-loss construction
Similar to VITS, we integrate Variational Autoencoder (VAE) and Generative Adversarial Network (GAN) methodologies into our training process. The overall loss is articulated as follows:
$L_{\text{total}}=L_{\text{recon}}+L_{\text{kl}}+L_{\text{adv}}(G)+L_{\text{fm}}(G)$ (1)
We transform the generated waveform to the mel-spectrogram domain and calculate the $L_{1}$ loss against the source speech’s mel-spectrogram to serve as the reconstruction loss:
$L_{\text{recon}}=\left|X_{\text{mel}}-\hat{X}_{\text{mel}}\right|$ (2)
Utilizing a linear-scale spectrogram $x_{\text{lin}}$, we acquire high-resolution information and optimize the distributions of our prior and posterior encoders via the Kullback-Leibler (KL) Loss:
$L_{\text{kl}}=\log q_{\phi}(z|x_{\text{lin}})-\log p_{\theta}(z|c)$ (3)
We apply a regularized flow $f_{\theta}$ over a factorized Gaussian prior distribution, enhancing the prior distribution through invertible transformations to elevate the authenticity and naturalness of the generated samples:
$q_{\phi}(z|x_{\text{lin}})=\mathcal{N}(z;\mu_{\phi}(x_{\text{lin}}),\sigma_{\phi}(x_{\text{lin}}))$ (4)
$p_{\theta}(z|c)=\mathcal{N}(f_{\theta}(z);\mu_{\theta}(c),\sigma_{\theta}(c))\left|\det\frac{\partial f_{\theta}(z)}{\partial z}\right|$ (5)
By introducing a pre-trained discriminator $D$, we ascertain the veracity and timbral consistency of the speech generated by the decoder $G$, leveraging $L_{\text{adv}}(G)$ for supervision. Furthermore, we incorporate an additional feature-matching Loss $L_{\text{fm}}(G)$ to ensure consistency in the reconstruction loss measured within the discriminator’s hidden layers between the generated and authentic speech.
II-F Training and testing strategy
During the training process, we input the same piece of speech into the network. The linear-scale spectrogram of the speech is sent to the Posterior Encoder to obtain the latent variable $Z$, which is then fed into the Decoder to produce synthesized speech. We calculate $L_{\text{recon}}$, $L_{\text{adv}}(G)$, and $L_{\text{fm}}(G)$. Simultaneously, under the constraints of multiple disentangled attributes, the latent variable $Z$ is transformed through a flow to $Z_{p}$, and $L_{\text{kl}}$ is computed with the content embedding derived from the Content Encoder.
During the inference process, we input the speech from two different speakers into the network. The content and emotion originate from the source speech, the timbre comes from the target speech, and the pitch is derived from a combined calculation of both the source and target speech. These elements serve as constraints for the flow, resulting in the latent variable $Z$, which is then fed into the decoder to produce synthesized speech.
III Experiment and Results
III-A Datasets
We conducted experiments on four datasets: VCTK, Opensinger, M4singer, and NUS-48E. We took the model weights that had been trained on the VCTK dataset (which contains 44 hours of speeches from 107 English speakers with various accents) as the starting point and then continued to train on the Opensinger dataset (a large-scale Chinese singing voice dataset) to learn the distinctive characteristics of Chinese singing voices. Finally, we used the M4singer dataset (a multi-style and multi-singer Chinese singing voice dataset) and the NUS-48E (an English singing voice dataset) to test the evaluation model’s capabilities in zero-shot singing voice conversion (SVC) and cross-domain conversion.
III-B detail
In the training stage, the speech resampling frequency is 16khz, and the speech Mel spectrogram is extracted by 512-point fast Fourier transform and 512-point window calculation. The model is preloaded with the pre-training weights of Quickvc[36] (trained on VCTK for 2 weeks), then trained on the Opensinger dataset for one and a half days, and tested on M4singer. At the same time, other experimental schemes such as FastSVC[37], SoVITS-Flow[38] are compared. All models have been fully iteratively trained on a single NVIDIA 3090ti, with a batch size of 48 and learning rates of 1e-4 and 5e-5 respectively.
III-C results and analysis
III-C1 Model evaluation
We adopt MOS/Naturalness, MOS/Similarity, Pesq, Voice Similarity and RTF as the evaluation indexes of the model, such as FastSVC, SoVITS-Flow [^1] https://github.com/svc-develop-team/so-vits-svc?tab=readme-ov-fle#sovits-model, CoMoSVC[38], DiffSVC[39], Multi-F0 Model[40], UCD-SVC[41],and the evaluation objects are the song conversion of different models in the dataset M4Singer → Opencpop and the cross-language song conversion. In Table 1, it shows that: (1) we evaluate the naturalness and similarity of voice by collecting people’s scores on MOS values extensively, and the results show that the naturalness of our proposed RASVC model reaches 4.14, which exceeds all baseline models except CoMoSVC, which proves that the multi-conditional fusion we joined has not reduced the naturalness, but made the emotion and pitch of singing more accurate and rich; (2) in terms of MOS similarity, people have scored more than the target timbre and the converted timbre. (3) Compared with other models, the 3)PESQ index of our proposed model has declined, which may be due to the uneven distribution of variables under most conditions, but it still remains at a high level. (4) We use a general speaker verification (ASV) model, such as ECAPA-TDNN,Comformer, to evaluate the timbre of the target song and the timbre of the converted song. The results show that our model maintains the highest sound similarity in cross-domain and cross-language singing conversion, which is closely related to F0 fundamental frequency assistance. (5) Finally, we tested the model RTF (the time required for the model to process one second of speech) in a single NVIDIA 3090ti. The other models here are from their source paper values, which may be because they use better hardware devices than us.
Figure 2: t-SNE visualization of transformed songs
III-C2 t-SNE visualization of transformed songs
As shown in Figure 2, we will transform the songs during the training process through the speaker encoder and draw them embedded in 2D space with t-SNE. It is found that different songs are still separated after conversion, which shows that our multi-conditional embedding does not destroy the identity information of speakers, but also effectively enhances the similarity, opens the distance information between different speakers and learns reasonable representation information.
III-C3 Visualization analysis of spectrum diagram
As can be seen from the Figure 2, when the male turns to the female, the resonance peak moves upward, and the interval of vertical stripes is obviously widened; When a woman turns to a man, the formant moves downward, and the interval of vertical stripes is obviously narrowed. At the same time, the main spectral features of the original voice (such as the overall energy distribution of the spectrum) are preserved after the two transformations. It can be seen that our model not only accurately adjusts the fundamental frequency (f0) in the transformation of male and female singing, but also effectively preserves the content of singing.
Figure 3: Visualization analysis of spectrum diagram
III-C4 Contrast visualization of speaker similarity
We let the human judges listen to three songs first, namely: the target singer’s song, the source singer’s song, and the transformed song. After that, we changed the songs sung by the target singer and the source singer and asked the judges to compare the similarities between the three songs and the source singer one by one, and then compare the similarities with the target singer. There are four kinds of ratings: same absolute certainty, same uncertainty, different uncertainty, and different absolute certainty. The results of the similarity comparison are shown in the Figure 4, and the evaluation results show that our model can transform the singing voice similar to the target singer.
Figure 4: visualization of speaker similarity
IV Conclusion
In this paper, we propose a high-fidelity flow-based model based on multi-decoupling feature constraints. This model uses timbre, pitch, content, and emotion to assist the flow model to complete the song synthesis and conversion, and it is also used in diversified singing scenes. The inverse Fourier transform is also applied to the decoder to improve the conversion efficiency. The experimental results show that the naturalness and similarity of the speech after the conversion of our proposed model are 4.14 and 4.02. Finally, we give a demo and release a Pytorch trainer for singing voice conversion to promote further research in this field.